OpenStudy (anonymous):

30.788 is 35.8% of what number is 86 the correct answer? what is 523% of 952 is 4978.96 the correct answer?

OpenStudy (callisto):

Answer to both question - Yes (i) number required = 30.788 / 35.8% = 86 (ii) value you need = 952 x 523% = 4978.96